OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -4x - y greater than or equal to -8 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (texaschic101):

-4x - y >= -8 -y >= 4x - 8 y <= -4x + 8 if it involves an equal sign,then it is a solid line. So that leaves us with B or D. Lets test some points...(0,0) -4(0) - y(0) >= -8 (this is false)....this means you shade the region that does not contain the points (0,0)....and when graphed, (0,0) falls in the shaded region, therefore, your answer is B.